OpenStudy (anonymous):

Integrate xX^2+2x+4)\√(x^2-4x) Help I've been trying forever

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is a bit hard to read isn't it?

OpenStudy (misty1212):

\[\int \frac{x^2+2x+4}{\sqrt{x^2-4x}}dx\]?